# Convergence Analysis of Natural Gradient Descent for Over-parameterized Physics-Informed Neural Networks

### v1

@@ 8.1. Proof of Lemma 7
$\frac{\partial s_{p}(\bm{w})}{\partial\bm{w}_{r}}$ $=\frac{a_{r}}{\sqrt{n_{1}m}}[\sigma^{{}^{\prime\prime}}(\bm{w}_{r}^{T}x_{p})w_{r0}\bm{x}_{p}+\sigma^{{}^{\prime}}(\bm{w}_{r}^{T}\bm{x}_{p})\begin{pmatrix}1\\ \bm{0}_{d}\end{pmatrix}-\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{p})\|\bm{w}_{r1}\|_{2}^{2}\bm{x}_{p}$
$\quad-2\sigma^{{}^{\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{p})\begin{pmatrix}0\\ \bm{w}_{r1}\end{pmatrix}]$
@@
(1) For $p\in[n_{1}]$, define the event
$A_{p,r}=\{\exists\bm{w}:\|\bm{w}-\bm{w}_{r}(0)\|_{2}\leq R,I\{\bm{w}^{T}\bm{x}_{p}\geq 0\}\neq I\{\bm{w}_{r}(0)^{T}\bm{x}_{p}\geq 0\}\}.$
Similar to (44), we can deduce that for any $p\in[n_{1}]$,
$P(A_{p,r})\leq\frac{2R}{\sqrt{2\pi}}.$
@@
$\leq\frac{1}{\sqrt{n_{1}m}}\left[R(\|\bm{w}_{r}(0)\|_{2}+1)+I\{A_{p,r}\}(\|\bm{w}_{r}(0)\|_{2}^{2}+1)\right],$
where the second inequality follows from the fact $\|\bm{w}-\bm{w}_{r}(0)\|_{2}<R\leq 1$ and the definition of $A_{p,r}$.
@@
$\lesssim\sum\limits_{r=1}^{m}\left(\sum\limits_{p=1}^{n_{1}}\frac{1}{n_{1}m}(R(\|\bm{w}_{r}(0)\|_{2}+1)+I\{A_{p,r}\}(\|\bm{w}_{r}(0)\|_{2}^{2}+1))^{2}+\sum\limits_{j=1}^{n_{2}}\frac{1}{n_{2}m}(R\|\bm{w}_{r}(0)\|_{2}+R)^{2}\right)$
$\lesssim\frac{R^{2}}{m}\sum\limits_{r=1}^{m}(\|\bm{w}_{r}(0)\|_{2}^{2}+1)+\frac{1}{n_{1}m}\sum\limits_{p=1}^{n_{1}}\sum\limits_{r=1}^{m}I\{A_{p,r}\}(\|\bm{w}_{r}(0)\|_{2}^{4}+1)$
@@
$\quad+\frac{1}{n_{1}m}\sum\limits_{p=1}^{n_{1}}\sum\limits_{r=1}^{m}I\{A_{p,r}\}\left(\|\bm{w}_{r}(0)\|_{2}^{4}I\{\|\bm{w}_{r}(0)\|_{2}^{2}\leq M\}+\|\bm{w}_{r}(0)\|_{2}^{4}I\{\|\bm{w}_{r}(0)\|_{2}^{2}>M\}+1\right)$
$\lesssim\frac{R^{2}}{m}\sum\limits_{r=1}^{m}(\|\bm{w}_{r}(0)\|_{2}^{2}+1)+\frac{M^{2}}{n_{1}m}\sum\limits_{p=1}^{n_{1}}\sum\limits_{r=1}^{m}I\{A_{p,r}\}+\frac{1}{m}\sum\limits_{r=1}^{m}\|\bm{w}_{r}(0)\|_{2}^{4}I\{\|\bm{w}_{r}(0)\|_{2}^{2}>M\},$
@@
(2) Note that when $\sigma$ satisfies the assumption 5, $\sigma^{{}^{\prime}},\sigma^{{}^{\prime\prime}}$ and $\sigma^{{}^{\prime\prime\prime}}$ are all Lipschitz continuous. Thus we can obtain that
@@
$\lesssim R^{2}\left[d^{2}+\frac{d^{2}}{\sqrt{m}}\sqrt{\log(\frac{1}{\delta})}+\frac{d^{2}}{m}(\log(\frac{1}{\delta}))^{2}\right],$
where the last inequality follows from the fact that $\|\|\bm{w}_{r}(0)\|_{2}^{4}\|_{\psi_{\frac{1}{2}}}\lesssim d^{2}$ and Lemma 8. ∎
8.2. Proof of Theorem 3
We prove Theorem 3 by induction. Our induction is the following condition for the hidden weights.
@@ 8.2. Proof of Theorem 3
Instead of inducing on the convergence rate of the empirical loss function, as shown in Condition 2, we perform induction on the movements of the hidden weights. Because with Condition 3, we can directly derive the following convergence rate of the empirical loss function.
Corollary 3.
If Condition 3 holds for $t=0,\cdots,k$ and $R^{{}^{\prime}}\leq R$ and $R^{{}^{\prime\prime}}\lesssim\sqrt{\lambda_{0}}$, then

### v3

@@ 10. Proof of Section 4
$\bm{H}(\bm{w})=\bm{D}^{T}\bm{D},\quad\bm{D}=\left[\frac{\partial s_{1}(\bm{w})}{\partial\bm{w}},\cdots,\frac{\partial s_{n_{1}}(\bm{w})}{\partial\bm{w}},\frac{\partial h_{1}(\bm{w})}{\partial\bm{w}},\cdots,\frac{\partial h_{n_{2}}(\bm{w})}{\partial\bm{w}}\right],$
and $\bm{H}^{\infty}=\mathbb{E}_{\bm{w}\sim\mathcal{N}(\bm{0},\bm{I})}\bm{H}(\bm{w})$.
We denote $\varphi(\bm{x};\bm{w})=\sigma^{{}^{\prime}}(\bm{w}^{T}\bm{x})w_{0}-\sigma^{{}^{\prime\prime}}(\bm{w}^{T}\bm{x})\|\bm{w}_{1}\|_{2}^{2}$, where $\bm{w}=(w_{0},\bm{w}_{1}^{T})^{T}$, $w_{0}\in\mathbb{R},\bm{w}_{1}\in\mathbb{R}^{d}$, then
$\frac{\partial s_{p}(\bm{w})}{\partial\bm{w}_{r}}=\frac{1}{\sqrt{n}_{1}}\frac{a_{r}}{\sqrt{m}}\frac{\partial\varphi(\bm{x}_{p};\bm{w}_{r})}{\partial\bm{w}_{r}}.$
Similarly, we denote $\psi(\bm{y};\bm{w})=\sigma(\bm{w}^{T}\bm{y})$, then
$\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}=\frac{1}{\sqrt{n}_{2}}\frac{a_{r}}{\sqrt{m}}\frac{\partial\psi(\bm{y}_{j},\bm{w}_{r})}{\partial\bm{w}_{r}}.$
With the notations, we can deduce that
$H_{p,j}^{\infty}=\left\{\begin{aligned} \frac{1}{n_{1}}\mathbb{E}_{\bm{w}\sim\mathcal{N}(\bm{0},\bm{I})}\left\langle\frac{\partial\varphi(\bm{x}_{p};\bm{w})}{\partial\bm{w}},\frac{\partial\varphi(\bm{x}_{j};\bm{w})}{\partial\bm{w}}\right\rangle&,&1\leq p\leq n_{1},1\leq j\leq n_{1},\\ \frac{1}{\sqrt{n_{1}n_{2}}}\mathbb{E}_{\bm{w}\sim\mathcal{N}(\bm{0},\bm{I})}\left\langle\frac{\partial\varphi(\bm{x}_{p};\bm{w})}{\partial\bm{w}},\frac{\partial\psi(\bm{y}_{j};\bm{w})}{\partial\bm{w}}\right\rangle&,&1\leq p\leq n_{1},n_{1}+1\leq j\leq n_{1}+n_{2},\\ \frac{1}{n_{2}}\mathbb{E}_{\bm{w}\sim\mathcal{N}(\bm{0},\bm{I})}\left\langle\frac{\partial\psi(\bm{y}_{p};\bm{w})}{\partial\bm{w}},\frac{\partial\psi(\bm{y}_{j};\bm{w})}{\partial\bm{w}}\right\rangle&,&n_{1}+1\leq p\leq n_{1}+n_{2},n_{1}+1\leq j\leq n_{1}+n_{2},\end{aligned}\right.$
where $H_{p,j}^{\infty}$ is the $(p,j)$-th entry of $\bm{H}^{\infty}$.
The proof of this lemma requires tools from functional analysis. Let $\mathcal{H}$ be a Hilbert space of integrable $(d+2)$-dimensional vector fields on $\mathbb{R}^{d+2}$, i.e., $f\in\mathcal{H}$ if $\mathbb{E}_{\bm{w}\sim\mathcal{N}(\bm{0},\bm{I})}[\|f(\bm{w})\|_{2}^{2}]<\infty$. The inner product for any two elements $f,g$ in $\mathcal{H}$ is $\mathbb{E}_{\bm{w}\sim\mathcal{N}(\bm{0},\bm{I})}[\langle f(\bm{w}),g(\bm{w})\rangle]$. Thus, proving $\bm{H}^{\infty}$ is strictly positive definite is equivalent to show that
$\frac{\partial\varphi(\bm{x}_{1};\bm{w})}{\partial\bm{w}},\cdots,\frac{\partial\varphi(\bm{x}_{n_{1}};\bm{w})}{\partial\bm{w}},\frac{\partial\psi(\bm{y}_{1};\bm{w})}{\partial\bm{w}},\cdots,\frac{\partial\psi(\bm{y}_{n_{2}};\bm{w})}{\partial\bm{w}}\in\mathcal{H}$
are linearly independent. Suppose that there are $\alpha_{1},\cdots,\alpha_{n_{1}},\beta_{1},\cdots,\beta_{n_{2}}\in\mathbb{R}$ such that
$\alpha_{1}\frac{\partial\varphi(\bm{x}_{1};\bm{w})}{\partial\bm{w}}+\cdots+\alpha_{n_{1}}\frac{\partial\varphi(\bm{x}_{n_{1}};\bm{w})}{\partial\bm{w}}+\beta_{1}\frac{\partial\psi(\bm{y}_{1};\bm{w})}{\partial\bm{w}}+\cdots+\beta_{n_{2}}\frac{\partial\psi(\bm{y}_{n_{2}};\bm{w})}{\partial\bm{w}}=0\ in\ \mathcal{H}.$
This implies that
$\alpha_{1}\frac{\partial\varphi(\bm{x}_{1};\bm{w})}{\partial\bm{w}}+\cdots+\alpha_{n_{1}}\frac{\partial\varphi(\bm{x}_{n_{1}};\bm{w})}{\partial\bm{w}}+\beta_{1}\frac{\partial\psi(\bm{y}_{1};\bm{w})}{\partial\bm{w}}+\cdots+\beta_{n_{2}}\frac{\partial\psi(\bm{y}_{n_{2}};\bm{w})}{\partial\bm{w}}=0$ (63)
holds for all $\bm{w}\in\mathbb{R}^{d+1}$, as $\sigma(\cdot)$ is smooth.
First, we compute the derivatives of $\varphi$ and $\psi$ with respect to $\bm{w}$. For the $k$-th derivative of $\psi(\bm{y};\bm{w})$ with respect to $\bm{w}$, we have
$\frac{\partial^{k}\psi(\bm{y};\bm{w})}{\partial\bm{w}^{k}}=\sigma^{(k)}(\bm{w}^{T}\bm{y})\bm{y}^{\otimes(k)},$
where $\otimes$ denotes the tensor product.
For $\varphi(\bm{x};\bm{w})$, let $\varphi_{0}(\bm{x};\bm{w})=\sigma^{{}^{\prime}}(\bm{w}^{T}\bm{x})w_{0}$ and $\varphi_{i}(\bm{x};\bm{w})=\sigma^{{}^{\prime\prime}}(\bm{w}^{T}\bm{x})w_{i}^{2}$ for $1\leq i\leq d$. Then
$\varphi(\bm{x};\bm{w})=\varphi_{0}(\bm{x};\bm{w})-\sum\limits_{i=1}^{d}\varphi_{i}(\bm{x};\bm{w}).$
For the $k$-th derivative of $\varphi_{0}(\bm{x};\bm{w})$ with respect to $\bm{w}$, analogous to the Leibniz rule for the $k$-th derivative of the product of two scalar functions, we have
$\frac{\partial^{k}\varphi_{0}(\bm{x};\bm{w})}{\partial\bm{w}^{k}}$ $=\sigma^{(k+1)}(\bm{w}^{T}\bm{x})w_{0}\bm{x}^{\otimes(k)}+\sum\limits_{i=1}^{k}\bm{x}^{\otimes(i-1)}\otimes\bm{e}_{0}\otimes\bm{x}^{\otimes(k-i)}\sigma^{(k)}(\bm{w}^{T}\bm{x})$ (64)
$=\sigma^{(k+1)}(\bm{w}^{T}\bm{x})w_{0}\bm{x}^{\otimes(k)}+\sigma^{(k)}(\bm{w}^{T}\bm{x})\sum\limits_{i=1}^{k}\bm{e}_{0}^{(i)}\otimes\bm{x}^{\otimes(k)},$
where $\bm{e}_{0}=(1,0,\cdots,0)^{T}\in\mathbb{R}^{d+1}$ and in the second equality, $\bm{e}_{0}^{(i)}$ denotes that $\bm{e}_{0}$ is placed at the $i$-th position.
Similarly, for $\varphi_{i}(\bm{x};\bm{w})$ where $1\leq i\leq d$, taking $i=1$ as example, we have
$\frac{\partial^{k}\varphi_{1}(\bm{x};\bm{w})}{\partial\bm{w}^{k}}$ $=\sigma^{(k+2)}(\bm{w}^{T}\bm{x})w_{1}^{2}\bm{x}^{\otimes(k)}+2kw_{1}\sigma^{(k+1)}(\bm{w}^{T}\bm{x})\sum\limits_{i=1}^{k}\bm{e}_{1}^{(i)}\otimes\bm{x}^{\otimes(k-1)}$ (65)
$\quad+k(k-1)\sigma^{(k)}(\bm{w}^{T}\bm{x})\sum\limits_{1\leq i<j\leq k}\bm{e}_{1}^{(i)}\otimes\bm{e}_{1}^{(j)}\otimes\bm{x}^{\otimes(k-2)},$
where $\bm{e}_{i}\in\mathbb{R}^{d+1}$ is a vector with the $(i+1)$-th component equal to $1$ and all other components equal to 00, and $\bm{e}_{1}^{(i)}$ indicates that $\bm{e}_{1}$ is placed at the $i$-th position.
By combining the derivatives of $\varphi_{0}(\bm{x};\bm{w}),\cdots,\varphi_{d}(\bm{x};\bm{w})$ from equations (64) and (65), we can compute the $k$-th derivative of $w_{0}\sigma^{{}^{\prime}}(\bm{w}^{T}\bm{x})-\sum\limits_{i=1}^{d}w_{i}^{2}\sigma^{{}^{\prime\prime}}(\bm{w}^{T}\bm{x})$ as follows:
$\frac{\partial^{k}\varphi(\bm{x};\bm{w})}{\partial\bm{w}^{k}}$ $=\frac{\partial^{k}\varphi_{0}(x;\bm{w})}{\partial\bm{w}^{k}}-\sum\limits_{i=1}^{d}\frac{\partial^{k}\varphi_{i}(\bm{x};\bm{w})}{\partial\bm{w}^{k}}$ (66)
$=w_{0}\sigma^{(k+1)}(\bm{w}^{T}\bm{x})\bm{x}^{\otimes(k)}+\sigma^{(k)}(\bm{w}^{T}\bm{x})\sum\limits_{i=1}^{k}\bm{e}_{0}^{(i)}\otimes\bm{x}^{\otimes(k)}$
$\quad-\sum\limits_{t=1}^{d}\left[w_{t}^{2}\sigma^{(k+2)}(\bm{w}^{T}\bm{x})\bm{x}^{\otimes(k)}+2kw_{t}\sigma^{(k+1)}(\bm{w}^{T}\bm{x})\sum\limits_{i=1}^{k}\bm{e}_{t}^{(i)}\otimes\bm{x}^{\otimes(k-1)}\right.$
$\quad\left.+k(k-1)\sigma^{(k)}(\bm{w}^{T}\bm{x})\sum\limits_{1\leq i<j\leq k}\bm{e}_{t}^{(i)}\otimes\bm{e}_{t}^{(j)}\otimes\bm{x}^{\otimes(k-2)}\right].$
Note that when any two points in $\{\bm{x}_{1},\cdots,\bm{x}_{n_{1}},\bm{y}_{1},\cdots,\bm{y}_{n_{2}}\}$ are non-parallel, the tensors
$\bm{x}_{1}^{\otimes(n_{1}+n_{2})},\cdots,\bm{x}_{n_{1}}^{\otimes(n_{1}+n_{2})},\bm{y}_{1}^{\otimes(n_{1}+n_{2})},\cdots,\bm{y}_{n_{2}}^{\otimes(n_{1}+n_{2})}$
are linearly independent (see Lemma G.6 in [11]). This observation motivates us to take the $(k-1)$-th derivative of both sides of equation (63) with respect to $\bm{w}$, yielding
$\alpha_{1}\frac{\partial^{k}\varphi(\bm{x}_{1};\bm{w})}{\partial\bm{w}^{k}}+\cdots+\alpha_{n_{1}}\frac{\partial^{k}\varphi(\bm{x}_{n_{1}};\bm{w})}{\partial\bm{w}^{k}}+\beta_{1}\frac{\partial^{k}\psi(\bm{y}_{1};\bm{w})}{\partial\bm{w}^{k}}+\cdots+\beta_{n_{2}}\frac{\partial^{k}\psi(\bm{y}_{n_{2}};\bm{w})}{\partial\bm{w}^{k}}=0.$ (67)
Since this equation holds for all $\bm{w}\in\mathbb{R}^{d+1}$, we specifically consider $\bm{w}=(w_{0},\bm{0}_{d})$, where $w_{0}$ is to be determined. Under this condition, equation (67) becomes
$w_{0}\sum\limits_{p=1}^{n_{1}}\alpha_{p}\left[\sigma^{(k+1)}(w_{0}x_{p}^{0})\bm{x}_{p}^{\otimes(k)}\right]+\sum\limits_{p=1}^{n_{1}}\alpha_{p}\left[\sigma^{(k)}(w_{0}x_{p}^{0})\bm{z}_{p}\right]+\sum\limits_{j=1}^{n_{2}}\beta_{j}\left[\sigma^{(k)}(w_{0}y_{j}^{0})\bm{y}_{j}^{\otimes(k)}\right]=0,$ (68)
where the tensor $\bm{z}_{p}$ is defined as
$\bm{z}_{p}=\sum\limits_{i=1}^{k}\bm{e}_{0}^{(i)}\otimes\bm{x}_{p}^{\otimes(k)}-k(k-1)\sum\limits_{t=1}^{d}\sum\limits_{1\leq i<j\leq k}\bm{e}_{t}^{(i)}\otimes\bm{e}_{t}^{(j)}\otimes\bm{x}_{p}^{\otimes(k-2)}.$ (69)
By assumption, for any positive integer $n\geq 0$ we have $\lim\limits_{x\to+\infty}\frac{\sigma^{(n)}(x)}{\phi(x)}=c_{n}\neq 0$. We first consider the case where all input components satisfy $x_{1}^{0}=\cdots=x_{n_{1}}^{0}=y_{1}^{0}=\cdots=y_{n_{2}}^{0}=a>0$. Under this condition, equation (68) simplifies to:
$w_{0}\sigma^{(k+1)}(w_{0}a)\left[\sum\limits_{p=1}^{n_{1}}\alpha_{p}\bm{x}_{p}^{\otimes(k)}\right]+\sigma^{(k)}(w_{0}a)\left[\sum\limits_{p=1}^{n_{1}}\alpha_{p}\bm{z}_{p}\right]+\sigma^{(k)}(w_{0}a)\left[\sum\limits_{j=1}^{n_{2}}\beta_{j}\bm{y}_{j}^{\otimes(k)}\right]=0.$ (70)
Dividing both sides of equation (70) by $\phi(w_{0}a)$ yields
$w_{0}\frac{\sigma^{(k+1)}(w_{0}a)}{\phi(w_{0}a)}\left[\sum\limits_{p=1}^{n_{1}}\alpha_{p}\bm{x}_{p}^{\otimes(k)}\right]+\frac{\sigma^{(k)}(w_{0}a)}{\phi(w_{0}a)}\left[\sum\limits_{p=1}^{n_{1}}\alpha_{p}\bm{z}_{p}\right]+\frac{\sigma^{(k)}(w_{0}a)}{\phi(w_{0}a)}\left[\sum\limits_{j=1}^{n_{2}}\beta_{j}\bm{y}_{j}^{\otimes(k)}\right]=0.$ (71)
Now taking the limit as $w_{0}$ tends to positive infinity, we observe that $\frac{\sigma^{(k)}(w_{0}a)}{\phi(w_{0}a)}$ converges to a non-zero constant, while $w_{0}\frac{\sigma^{(k+1)}(w_{0}a)}{\phi(w_{0}a)}$ diverges to infinity. This asymptotic behavior leads to the following conclusions:
$\sum\limits_{p=1}^{n_{1}}\alpha_{p}\bm{x}_{p}^{\otimes(k)}=0,\ \sum\limits_{p=1}^{n_{1}}\alpha_{p}\bm{z}_{p}+\sum\limits_{j=1}^{n_{2}}\beta_{j}\bm{y}_{j}^{\otimes(k)}=0.$ (72)
By the linear independence of the tensor products (established earlier), we can deduce that $\alpha_{p}=0$ for all $p=1,\cdots,n_{1}$, which subsequently implies $\sum\limits_{j=1}^{n_{2}}\beta_{j}\bm{y}_{j}^{\otimes(k)}=0$ and thus $\beta_{j}=0$ for all $j=1,\cdots,n_{2}$.
When our previous assumption does not hold—that is, when $x_{1}^{0},\cdots,x_{n_{1}}^{0},y_{1}^{0},\cdots,y_{n_{2}}^{0}$ are not necessarily all equal—we proceed as follows:
Case 1: All elements are strictly positive.
Let $b=\min\{x_{1}^{0},\cdots,x_{n_{1}}^{0},y_{1}^{0},\cdots,y_{n_{2}}^{0}\}$. Dividing both sides of (68) by $\phi(w_{0}b)$, we observe that for any $x>b$,
$\lim\limits_{w_{0}\to+\infty}\frac{\sigma^{(n)}(w_{0}x)}{\phi(w_{0}b)}=0.$
Thus, the problem reduces to the previously considered case where all inputs are equal (to $b$). Due to linear independence, the coefficients $\alpha_{p}$ and $\beta_{j}$ corresponding to the minimal $b$ must vanish. Repeating this process iteratively, we conclude that all $\alpha_{p}$ and $\beta_{j}$ must be zero.
Case 2: Some elements are zero.
Since the $\bm{x}_{p}$ are interior points (and thus non-zero), any zero-valued inputs must correspond to boundary or initial conditions. Let us assume without loss of generality that $y_{1}^{0},\cdots,y_{n_{2}}^{0}$ are all zero.
1. If $\sigma^{(k)}(0)=0$, , then following our previous method, we can conclude that the coefficients corresponding to non-zero inputs vanish. Returning to equation (67), we have
$\sum\limits_{j=1}^{n_{2}}\beta_{j}\sigma^{(k)}(\bm{w}^{T}\bm{y}_{j})\bm{y}_{j}^{\otimes(k)}=0.$ (73)
By the independence of $\bm{y}_{1},\cdots,\bm{y}_{n_{2}}$, we can deduce that $\beta_{j}\sigma^{(k)}(\bm{w}^{T}\bm{y}_{j})=0$ holds for all $j\in[n_{2}]$. From the assumption, we can select $\bm{w}$ such that $\sigma^{(k)}(\bm{w}^{T}\bm{y}_{j})\neq 0$, and consequently, $\beta_{j}=0$ holds for all $j\in[n_{2}]$.
2. If $\sigma^{(k)}(0)\neq 0$, let $b$ be the smallest strictly positive value among $x_{1}^{0},\cdots,x_{n_{1}}^{0},y_{1}^{0},\cdots,y_{n_{2}}^{0}$. Divide (70) by $\phi(w_{0}b/2)$. Since all other positive terms decay to zero as $w_{0}\to+\infty$, we obtain:
$\lim\limits_{w_{0}\to\infty}\sum\limits_{j=1}^{n_{2}}\beta_{j}\sigma^{(k)}(0)\bm{y}_{j}^{\otimes(k)}/\phi(w_{0}b/2)=0.$ (74)
This implies that $\sum\limits_{j=1}^{n_{2}}\beta_{j}\bm{y}_{j}^{\otimes(k)}=0$. By linear independence, all $\beta_{j}=0$. ∎
10.2. Proof of Lemma 4.5
Proof.
Recall that
@@
$\frac{\partial s_{p}(\bm{w})}{\partial\bm{w}_{r}}$ $=\frac{a_{r}}{\sqrt{n_{1}m}}\left[\sigma^{{}^{\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{p})w_{r0}\bm{x}_{p}+\sigma^{{}^{\prime}}(\bm{w}_{r}^{T}\bm{x}_{p})\begin{pmatrix}1\\ \bm{0}_{d+1}\end{pmatrix}-\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{p})\|\bm{w}_{r1}\|_{2}^{2}\bm{x}_{p}\right.$
$\quad\left.-2\sigma^{{}^{\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{p})\begin{pmatrix}0\\ \bm{w}_{r1}\end{pmatrix}\right]$
@@
(1) When $\sigma(\cdot)$ is the $\text{ReLU}^{3}$ activation function.
@@
$\leq\frac{1}{\sqrt{n_{1}m}}\left[R(\|\bm{w}_{r}(0)\|_{2}+1)+I\{A_{pr}\}(\|\bm{w}_{r}(0)\|_{2}^{2}+1)\right],$
where the second inequality follows from the fact $\|\bm{w}-\bm{w}_{r}(0)\|_{2}<R\leq 1$ and the definition of $A_{pr}$ in (26).
@@
$\lesssim\sum\limits_{r=1}^{m}\left(\sum\limits_{p=1}^{n_{1}}\frac{1}{n_{1}m}\left(R(\|\bm{w}_{r}(0)\|_{2}+1)+I\{A_{pr}\}(\|\bm{w}_{r}(0)\|_{2}^{2}+1)\right)^{2}+\sum\limits_{j=1}^{n_{2}}\frac{1}{n_{2}m}(R\|\bm{w}_{r}(0)\|_{2}+R)^{2}\right)$
$\lesssim\frac{R^{2}}{m}\sum\limits_{r=1}^{m}(\|\bm{w}_{r}(0)\|_{2}^{2}+1)+\frac{1}{n_{1}m}\sum\limits_{p=1}^{n_{1}}\sum\limits_{r=1}^{m}I\{A_{pr}\}(\|\bm{w}_{r}(0)\|_{2}^{4}+1)$
@@
$\quad+\frac{1}{n_{1}m}\sum\limits_{p=1}^{n_{1}}\sum\limits_{r=1}^{m}I\{A_{pr}\}\left(\|\bm{w}_{r}(0)\|_{2}^{4}I\{\|\bm{w}_{r}(0)\|_{2}^{2}\leq M\}+\|\bm{w}_{r}(0)\|_{2}^{4}I\{\|\bm{w}_{r}(0)\|_{2}^{2}>M\}+1\right)$
$\lesssim\frac{R^{2}}{m}\sum\limits_{r=1}^{m}(\|\bm{w}_{r}(0)\|_{2}^{2}+1)+\frac{M^{2}}{n_{1}m}\sum\limits_{p=1}^{n_{1}}\sum\limits_{r=1}^{m}I\{A_{pr}\}+\frac{1}{m}\sum\limits_{r=1}^{m}\|\bm{w}_{r}(0)\|_{2}^{4}I\{\|\bm{w}_{r}(0)\|_{2}^{2}>M\},$
@@
(2) Note that when $\sigma$ satisfies Assumption 4.3, $\sigma^{{}^{\prime}},\sigma^{{}^{\prime\prime}}$ and $\sigma^{{}^{\prime\prime\prime}}$ are all Lipschitz continuous and bounded. Thus, we can obtain that
@@
$\lesssim R^{2}\left[d^{2}+\frac{d^{2}}{\sqrt{m}}\sqrt{\log\left(\frac{1}{\delta}\right)}+\frac{d^{2}}{m}\left(\log\left(\frac{1}{\delta}\right)\right)^{2}\right],$
where the last inequality follows from the fact that $\left\|\|\bm{w}_{r}(0)\|_{2}^{4}\right\|_{\psi_{\frac{1}{2}}}\lesssim d^{2}$ and Lemma D.4. ∎
10.3. Proof of Theorem 4.7
For the sake of completeness in the proof, we restate Condition 1 and Corollary 4.11 from the main text, and label them as Condition 3 and Corollary C.1, respectively.
@@
Corollary 10.1.
If Condition 3 holds for $t=0,\cdots,k$ and $R^{{}^{\prime}}\leq R$ and $R^{{}^{\prime\prime}}\lesssim\sqrt{1-\eta}\sqrt{\lambda_{0}}$, then